OpenStudy (anonymous):

FAN AND MEDAL FOR STEPS AND EXPLAINNING !!! Solve using Quadratic Formula. 3x^2-5=0

OpenStudy (anonymous):

+5 both sides, get??

OpenStudy (anonymous):

3x^2=5 ??

OpenStudy (anonymous):

Do you know what the quadratic formula is for a polynomial of the form \(ax^2+bx+c\) ?

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

divide both sides by 3 , get?

OpenStudy (anonymous):

x^2 = 5 over 3 ??

OpenStudy (anonymous):

Just some background information: If you have some polynomial of the form \(ax^2+bx+c\), The quadratic formula gives you the values of \(x\) that, when plugged into the polynomial, give you the value of \(0\).

OpenStudy (anonymous):

square root both sides, get?

OpenStudy (anonymous):

OOOPS, the problem says to solve the equation using the quadratic formula, not manipulation.

OpenStudy (anonymous):

3x^2+0-5=0 ??

OpenStudy (anonymous):

oh yea, I am sorry. just take the easiest way to get the answer. I am SO SORRY

OpenStudy (anonymous):

its ok

OpenStudy (anonymous):

@OOOPS, if the purpose of the problem is to practice the quadratic formula, then he should practice the quadratic formula. And obeydallas13, yes, that is correct. Now apply the quadratic formula to that polynomial.

OpenStudy (anonymous):

\[x=-0\pm \sqrt{0^2-4(3)(5)}\]

OpenStudy (anonymous):

Very good so far, and very close. But does \(c=5\) or \(-5\)?

OpenStudy (anonymous):

-5 sorry

OpenStudy (anonymous):

Also, you are forgetting to divide the whole thing by something. Now try to redo it. You are very close :).

OpenStudy (anonymous):

\[x=-\pm \sqrt{0^2-4(3)(-5)} \over 2(3)\]

OpenStudy (anonymous):

Perfect, and I assume you meant \(-0\) before that \(\pm\) sign. Now which values of \(x\) do you get with this?

OpenStudy (anonymous):

yea I meant that and what do you mean of x

OpenStudy (anonymous):

Solving the quadratic formula gives you values of \(x\). If you plug these values of \(x\) into the polynomial, it will equal \(0\). Basically I just mean: solve what you got.

OpenStudy (anonymous):

\[x=\pm \sqrt{60} \over 6\]

OpenStudy (anonymous):

Yep. Does the problem say to simplify? Also you don't mean to have that "\(x=\)" on that fraction right?

OpenStudy (anonymous):

You meant \(x=\frac{\pm\sqrt{60}}{6}\) right?

OpenStudy (anonymous):

\[\pm \sqrt{60} \over 6\] ???

OpenStudy (anonymous):

\[x=\frac{ \pm60 }{ 6 }\]

OpenStudy (anonymous):

Yes. You got the correct answer the first time. FYI: The significance of these values is that if you plug in \(\frac{\sqrt{60}}{6}\) or \(-\frac{\sqrt{60}}{6}\) into the quadratic polynomial, it will equal zero.

OpenStudy (anonymous):

thank you for your help !!! I think your expecting something right ??

OpenStudy (anonymous):

Lol I don't do this for the medals, I just hope that you found me helpful :). Also, remember that the quadratic equation ONLY works if you want the quadratic to equal \(0\), nothing else.

OpenStudy (anonymous):

And feel free to tag/message me or something if you need help in the future. Bye bye!